# Data Valuation using Neural Networks for Efficient Instruction Fine-Tuning

### v2

@@ 5 Subset Selection Evaluation
We use MixInstruct as well as Alpaca (Taori et al., 2023) to evaluate NN-CIFT. These are both instruction tuning datasets where we use 15k for training, 5k for validation, and 5k for testing. We evaluate using two models: microsoft/Phi-3-small-8k-instruct (Abdin et al., 2024) and meta-llama/Llama-3.1-8B (Grattafiori et al., 2024). Note: we use Phi-3 and Llama-8B as shorthand for these models, respectively. Phi-3 has 7.39B parameters and LLama-8B has 8.03B parameters.
@@
To evaluate the instruction following capabilities of our fine-tuned model $\mathcal{M}^{\prime}$, we employ a variety of metrics to capture the similarity between ground truth answers and predicted answers from $\mathcal{M}^{\prime}$: (1) ROUGE (Lin, 2004): $n$-gram word overlap (specifically, rouge-1), (2) BGE: semantic similarity of embeddings using bge-large-en-v1.5, and (3) LAJ: an LLM-as-a-Judge, namely the prometheus-7b-v2.0 model (Kim et al., 2023). Prometheus’ grading rubric is borrowed from Agarwal et al. (2025) in Appendix B. Next, to evaluate the costs of each method, we use time (in seconds) took on 2 Nvidia A40 GPUs.
@@
We use $u=0.05$ for training the InfluenceNetwork. We also use a small fraction of $\mathcal{D_{F}}$ to fine-tune the language model – we call this fraction $v$. We evaluate with $v=0.3$. Our evaluation framework includes two different settings to fine-tune the language model: using the selected subset of data points as (1) PEFT data for QLoRA (Dettmers et al., 2023) on $\mathcal{M}$, or (2) in-context learning (ICL) examples. To elaborate on the ICL set up, we choose the top-5 most semantically similar samples from the chosen subset to add in-context. To measure semantic similarity, we again use bge-large-en-v1.5. Table 2 reports results for Phi-3 on both datasets with $v=0.3$; Table 3 reports results for Llama-8B on both datasets with $v=0.3$; Table 4 reports the cost in time for each method. All tables report the results for one run.
@@
Table 4: Costs (in seconds) of data valuation. Following are the specifications on each method. Random: choosing a random subset of points as a subset. SelectIT: calculating the ranking scores for each data point according to Appendix B.2. LESS: computing the cosine similarity between pairs of projected gradients for $\mathcal{D_{F}}$ and $\mathcal{D_{T}}$, according to Equation 3. DELIFT (SE): computing the distance between each pair of embeddings $(i,j):i\in\mathcal{D_{F}},j\in\mathcal{D_{T}}$, according to Equation 2. DELIFT: computing the inference-based utility metric for each pair of embeddings $(i,j)$, according to Equation 1. NN-CIFT: Steps 1 and 2 in Figure 1. Note, the costs of DistilGPT2 are the same across both models because they use the same data valuation (Phi-3 or Llama-8B are used for data selection/evaluation).
5.1 Analysis
Table 4 reports the costs for each method, in seconds. It shows that data valuation can be performed at 77-99% faster than the original influence functions. This is because the number of parameters in NN-CIFT is 0.0026-0.0029% the size of the language model in the original influence function. Also, when using the DistilGPT2 model, which is near 1% the size of the language model, the costs are reduced by 54-91%. While these results are promising, the results on the downstream task of subset selection clearly differentiate NN-CIFT and the DistilGPT2 baseline. To elaborate, despite the significant speedups, NN-CIFT shows no compromise to performance. Table 2 reports the results for Phi-3 and Table 3 reports the results for Llama-8B for $v=0.3$.
Keeping these in mind, NN-CIFT is able to achieve performance comparable to the original data valuation methods, even across models and datasets. However, DistilGPT2 shows performance degradations, which is more pronounced in the model-dependent methods (DELIFT, LESS, and SelectIT). This is because the model-dependent methods experience significant performance gains when the data valuation model is the same as the fine-tuning model.
The absolute average performance difference across metrics between the original influence functions and NN-CIFT is only 1.40% [^2] The average performance difference is calculated by taking the absolute difference in performance, dividing it by the original performance, and then averaging this ratio across all settings (datasets, methods, metrics, baselines).. Because the neural network is able to estimate the influence values with great accuracy, the selected subsets of data would be mostly the same between the original influence function and NN-CIFT. Hence, the performance difference of 1.40% can be attributed as the variability in the language model’s performance between two runs. Additionally, this trend is consistent across datasets and models, which shows the wide applicability of our method.
Figure 4: Hyperparameter study for $u$ and $v$ on MixInstruct with DELIFT’s influence function. Lighter colors indicate better BGE performance.
5.2 Hyperparameter study #2: Trade-off between $u$ and $v$
@@
As expected, we notice a few trends. (1) QLoRA generally has better performance than ICL. This is because fine-tuning has more impact on the model than simply adding examples to the prompt (i.e., prompt engineering). (2) The bottom right tends to be darker as fewer IFT data lead to insufficient training. (3) Larger IFT subsets, especially in the ICL setting, lead to poorer performance. During ICL, the top-5 semantically similar samples are chosen from the subset to add as in-context examples. However, semantic similarity does not always translate to performance enhancement as these samples can be harmful to the model’s performance. Finally, a follow-up to (3), the highest performance regions tend to be around $v=0.2$ $0.4$. Appendix A contains results on smaller subsets of IFT data ($v=0.1$ and $0.2$).
6 Conclusion
In this paper, we introduce NN-CIFT: Neural Networks for effiCient Instruction Fine-Tuning to distill highly parameterized models used in modern influence functions into small neural networks. We empirically show the effectiveness of our InfluenceNetwork design through low prediction error rates, and competitive performance on the downstream task of subset selection for IFT. We use four different influence functions to test with NN-CIFT; our experimentation shows that NN-CIFT can lower costs for expensive data valuation, is adaptive to all kinds of influence functions (model-dependent or -independent; pairwise or pointwise), and does not require retraining for new data.
7 Limitations
While NN-CIFT is effective, it is heavily dependent on the influence function. The influence functions that were studied require large datasets to be annotated, which can be infeasible. Furthermore, NN-CIFT cannot yet be used for areas such as task-specific dataset selection or continual learning. In these cases, the objectives of data selection are beyond representation. Finally, even though costs were shown to be much smaller, NN-CIFT still incurs a quadratic cost. Although we show results for SelectIT, which runs in linear time, SelectIT is not able to outperform the pairwise methods. Future work will involve finding a solution that can estimate influence of a data point to a data set (or a model’s training dynamics) while also incurring linear time cost.
Appendix A Evaluation on smaller subsets
Tables 5 and 7 report extra results for the Phi-3 model on $v=0.1$ and $v=0.2$, respectively. Similarly, Tables 6 and 8 report results for Llama-8B on $v=0.1$ and $v=0.2$, respectively. With Tables 2 and 3 in the main text, these results show an increasing trend in performance with a higher subset of IFT data (i.e., higher $v$). They also show similar trends where NN-CIFT performs similarly to the original influence function.
@@ Appendix A Evaluation on smaller subsets
Table 7: Results on the Llama-8b model with $v=0.2,u=0.05$. NN-CIFT + Method and DistilGPT2 + Method follow the same definitions as in Table 2. The average performance difference between NN-CIFT and the original influence function is merely 1.08%.
@@
Table 8: Results on the Llama-8b model with $v=0.2,u=0.05$. NN-CIFT + Method and DistilGPT2 + Method follow the same definitions as in Table 2. The average performance difference between NN-CIFT and the original influence function is merely 1.26%.
Appendix B Influence Functions
@@
B.1 Pairwise Influence Functions
@@
(Agarwal et al., 2025) is a model-dependent, inference-based metric. Samples $(i_{x},i_{y})\in\mathcal{D_{F}}$ are used as in-context examples for evaluating $(j_{x},j_{y})\in\mathcal{D_{T}}$, and those with improved model performance are chosen to represent $\mathcal{D_{T}}$. This can be calculated by comparing the performance with and without $(i_{x},i_{y})$ as an in-context example (where $D(\cdot,\cdot)\in[0,1]$ is a function to measure distance between two probability distributions, and $f(q|\theta)$ is a language model with parameters $\theta$ and input query $q$):
@@
After data valuation, the data selection stage consists of using submodular functions (Bilmes, 2022). In particular, we use the Facility Location submodular function. It takes as input a similarity kernel that will optimize the maximum similarity between the chosen subset and the overall dataset while also minimizing the size of the chosen subset. To minimize the subset size, the Facility Location – and submodular functions, in general – employ a diminishing gains property. This property states that samples added to a smaller subset have more value than samples added to a larger subset. Hence, we rely on our influence function to capture the informativeness of samples, and submodular functions to choose a set of representative samples, resulting in a small, information-rich subset on which to fine-tune a model.
@@
(Agarwal et al., 2025) is a model-independent metric, and chooses samples from $\mathcal{D_{F}}$ which are semantically closest to the samples from $\mathcal{D_{T}}$. Semantic distance is calculated by the cosine distance between embeddings of samples:
$\text{sim}(i,j)=\frac{<\text{emb}((i_{x},i_{y})),\text{emb}((j_{x},j_{y}))>}{||\text{emb}((i_{x},i_{y}))||\cdot||\text{emb}((j_{x},j_{y}))||}$ (2)
@@
(Xia et al., 2024) is model-dependent, gradient-based metric. Here, gradients between samples in $\mathcal{D_{F}}$ and $\mathcal{D_{T}}$ are matched by cosine similarity, and those that match the highest are chosen to represent $\mathcal{D_{T}}$ (where $\nabla(q;\theta)$ is the gradient of data point $q$ from a model with parameters $\theta$):
$\text{sim}(i,j)=\frac{<\nabla((i_{x},i_{y});\theta),\nabla((j_{x},j_{y});\theta)>}{||\nabla((i_{x},i_{y});\theta)||\cdot||\nabla((j_{x},j_{y});\theta)||}$ (3)
@@
B.2 Pointwise Influence Functions
Finally, SelectIT Liu et al. (2024a) is another model-dependent metric that uses performance signals for data valuation, but incurs linear cost as it uses a model’s uncertainty to rank data samples. Still, as mentioned in Table 1 from the main text, the linear time operations are forward propagations through LLMs.

### v4

@@ 5 Subset Selection Evaluation
We use MixInstruct, (Taori et al., 2023), Alpaca (Taori et al., 2023), and MMLU (Hendrycks et al., 2021) to evaluate NN-CIFT. These are instruction-tuning, preference alignment, and knowledge-based benchmarks where we use 15k for training, 5k for validation, and 5k for testing. We evaluate using three models: meta-llama/Llama-3.1-8B (Grattafiori et al., 2024), Qwen/Qwen2.5-1.5B (Qwen et al., 2025), and mistralai/Mistral-Small-Instruct-2409 (22.2B) (Jiang et al., ). Note, we use Llama-8b, Qwen2.5 and Mistral as shorthand for the rest of the experimental section.
@@
To evaluate the instruction following capabilities of our fine-tuned model $\mathcal{M}^{\prime}$, we employ a variety of metrics to capture the similarity between ground truth answers and predicted answers from $\mathcal{M}^{\prime}$: (1) ROUGE (Lin, 2004): $n$-gram word overlap (specifically, rouge-1), (2) BGE: semantic similarity of embeddings using bge-large-en-v1.5, and (3) LAJ: an LLM-as-a-Judge, namely the prometheus-7b-v2.0 model (Kim et al., 2023). Prometheus’ grading rubric is borrowed from Agarwal et al. (2025) in Appendix B. Next, to evaluate the costs of each method, we use time (in seconds) took on 2 Nvidia A40 GPUs. However, for MMLU, we use classification accuracy.
@@
We use $u=0.05$ for training the InfluenceNetwork. We also use a small fraction of $\mathcal{D_{F}}$ to fine-tune the language model – we call this fraction $v$. We evaluate with $v=0.3$. Our evaluation framework includes two different settings to fine-tune the language model: using the selected subset of data points as (1) PEFT data for QLoRA (Dettmers et al., 2023) on $\mathcal{M}$, or (2) in-context learning (ICL) examples. To elaborate on the ICL set up, we choose the top-5 most semantically similar samples from the chosen subset to add in-context. To measure semantic similarity, we again use bge-large-en-v1.5. Table 4-6 reports results for each model on all three datasets with $v=0.3$; Table 7 reports the cost in time for each method. All tables report the results for one run.
5.1 Analysis
Table 7 reports the costs for each method, in seconds. It shows that data valuation can be performed at 77-99% faster than the original influence functions. This is because the number of parameters in NN-CIFT is 0.00096-0.013% the size of the language model in the original influence function. Also, when using the DistilGPT2 model, which is near 1% the size of the language model, the costs are reduced by 54-91%. While these results are promising, the results on the downstream task of subset selection clearly differentiate NN-CIFT and the DistilGPT2 baseline. Despite the significant speedups, NN-CIFT shows no compromise to performance, as shown in Tables 4-6.
Keeping these in mind, NN-CIFT is able to achieve performance comparable to the original data valuation methods, even across models and datasets. DistilGPT2 shows performance degradations, especially in the model-dependent methods (DELIFT, LESS, and SelectIT). This is because the model-dependent methods experience significant performance gains when the data valuation model is the same as the fine-tuning model. We note that our evaluation’s focus is that NN-CIFT works as well as the original influence functions, and not the comparison of performance between them.
The absolute average performance difference across metrics between the original influence functions and NN-CIFT is only 0.16% [^2] The average performance difference is calculated by taking the absolute difference in performance, dividing it by the original performance, and then averaging this ratio across all settings (datasets, methods, metrics, baselines).. Because the neural network is able to estimate the influence values with great accuracy, the selected subsets of data would be mostly the same between the original influence function and NN-CIFT. Hence, the performance difference of 0.16% can be attributed as the variability in the language model’s performance between two runs. Additionally, this trend is consistent across datasets and models, which shows the wide applicability of our method.
@@ 5.1 Analysis
Table 7: Costs (in seconds) of data valuation. Specifically: Random: chooses a random subset of points. SelectIT: calculates the ranking scores for each data point according to Appendix C.2. LESS: computes the cosine similarity between pairs of projected gradients for $\mathcal{D_{F}}$ and $\mathcal{D_{T}}$, according to Equation 6. DELIFT (SE): computes the distance between each pair of embeddings $(i,j):i\in\mathcal{D_{F}},j\in\mathcal{D_{T}}$, according to Equation 5. DELIFT: computes the inference-based utility metric for each pair of embeddings $(i,j)$, according to Equation 4. NN-CIFT: Steps 1 and 2 in Figure 1. Note, the costs of DistilGPT2 are the same across both models because they use the same data valuation.
6 Conclusion
In this paper, we introduce NN-CIFT: Neural Networks for effiCient Instruction Fine-Tuning to distill highly parameterized models used in modern influence functions into small neural networks. We empirically show the effectiveness of our InfluenceNetwork design through low prediction error rates, and competitive performance on the downstream task of subset selection for IFT. We use four different influence functions to test with NN-CIFT; our experimentation shows that NN-CIFT can lower costs for expensive data valuation, is adaptive to all kinds of influence functions (model-dependent or -independent; pairwise or pointwise), and does not require retraining for new data. Future work will focus on two things: applicability and more targeted modeling. To improve applicability, incorporate more fine-tuning stage objectives such as task-specific dataset selection or continual learning. To target the modeling, we can shift from learning influence between data points to estimating influence between data and a model parameters. Finally, we leave to future work to generalize NN-CIFT to distributional shifts within the data (although, the training is lightweight enough that retraining would suffice for now).
Appendix A Hyperparameter Studies
A.1 Hyperparameter Study #1: InfluenceNetwork sizes
A.2 Hyperparameter study #2: Trade-off between $u$ and $v$
@@
Table 8: Hyperparameter study for $u$ and $v$ on MixInstruct with DELIFT’s influence function. Lighter colors indicate better BGE performance.
As expected, we notice a few trends. (1) QLoRA generally has better performance than ICL. This is because fine-tuning has more impact on the model than simply adding examples to the prompt (i.e., prompt engineering). (2) The bottom right tends to be darker as fewer IFT data lead to insufficient training. (3) Larger IFT subsets, especially in the ICL setting, lead to poorer performance. During ICL, the top-5 semantically similar samples are chosen from the subset to add as in-context examples. However, semantic similarity does not always translate to performance enhancement as these samples can be harmful to the model’s performance. Finally, a follow-up to (3), the highest performance regions tend to be around $v=0.2$ - $0.4$. Appendix B contains results on smaller subsets of IFT data ($v=0.1$ and $0.2$).
Appendix B Evaluation on Smaller Subsets
Tables 9 and 11 report extra results for the Phi-3 model on $v=0.1$ and $v=0.2$, respectively. Similarly, Tables 10 and 12 report results for Llama-8B on $v=0.1$ and $v=0.2$, respectively. With Table 4 in the main text, these results show an increasing trend in performance with a higher subset of IFT data (i.e., higher $v$). They also show similar trends where NN-CIFT performs similarly to the original influence function.
@@ Appendix B Evaluation on Smaller Subsets
Table 11: Results on the Llama-8b model with $v=0.2,u=0.05$. NN-CIFT + Method and DistilGPT2 + Method follow the same definitions as in Table LABEL:table:_phi_v=0.3. The average performance difference between NN-CIFT and the original influence function is merely 1.08%.
@@
Table 12: Results on the Llama-8b model with $v=0.2,u=0.05$. NN-CIFT + Method and DistilGPT2 + Method follow the same definitions as in Table LABEL:table:_phi_v=0.3. The average performance difference between NN-CIFT and the original influence function is merely 1.26%.
Appendix C Influence Functions
@@
C.1 Pairwise Influence Functions
@@
[Agarwal et al., 2025] is a model-dependent, inference-based metric. Samples $(i_{x},i_{y})\in\mathcal{D_{F}}$ are used as in-context examples for evaluating $(j_{x},j_{y})\in\mathcal{D_{T}}$, and those with improved model performance are chosen to represent $\mathcal{D_{T}}$. This can be calculated by comparing the performance with and without $(i_{x},i_{y})$ as an in-context example (where $D(\cdot,\cdot)\in[0,1]$ is a function to measure distance between two probability distributions, and $f(q|\theta)$ is a language model with parameters $\theta$ and input query $q$):
@@
After data valuation, the data selection stage consists of using submodular functions [Bilmes, 2022]. In particular, we use the Facility Location submodular function. It takes as input a similarity kernel that will optimize the maximum similarity between the chosen subset and the overall dataset while also minimizing the size of the chosen subset. To minimize the subset size, the Facility Location – and submodular functions, in general – employ a diminishing gains property. This property states that samples added to a smaller subset have more value than samples added to a larger subset. Hence, we rely on our influence function to capture the informativeness of samples, and submodular functions to choose a set of representative samples, resulting in a small, information-rich subset on which to fine-tune a model.
@@
[Agarwal et al., 2025] is a model-independent metric, and chooses samples from $\mathcal{D_{F}}$ which are semantically closest to the samples from $\mathcal{D_{T}}$. Semantic distance is calculated by the cosine distance between embeddings of samples:
[…]
@@
[Xia et al., 2024] is model-dependent, gradient-based metric. Here, gradients between samples in $\mathcal{D_{F}}$ and $\mathcal{D_{T}}$ are matched by cosine similarity, and those that match the highest are chosen to represent $\mathcal{D_{T}}$ (where $\nabla(q;\theta)$ is the gradient of data point $q$ from a model with parameters $\theta$):
[…]
@@
C.2 Pointwise Influence Functions
Finally, SelectIT Liu et al. [2024a] is another model-dependent metric that uses performance signals for data valuation, but incurs linear cost as it uses a model’s uncertainty to rank data samples. Still, as mentioned in Table 1 from the main text, the linear time operations are forward propagations through LLMs.